OpenStudy (anonymous):

what mixed number is equivalent to this decimal? 8.00027

Vocaloid (vocaloid):

any ideas on how to write 0.00027 as a fraction?

OpenStudy (anonymous):

nope toooo big

OpenStudy (anonymous):

also i like listening to vocaloid songs especially bio-hazard and repeat

Vocaloid (vocaloid):

well, let me give you a hint 27/10 = 2.7 27/100 =.27 27/1000 = 0.027 do you think you can figure out how to write 0.00027 as a fraction?

OpenStudy (anonymous):

nope

Vocaloid (vocaloid):

think carefully about what I wrote. follow the pattern

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

27/10000

OpenStudy (anonymous):

@Vocaloid

OpenStudy (mathmale):

Your result looks good. Now re-write 8.00027 as a mixed number.